I am being asked to setup a system of linear equations for the following word problem & then use either substitution or linear combinations to solve.: There were a total of 480 tickets sold for the school play. Adult tickets sold for $8 & student tickets sold for $5. the ticket sales brought in &3,369. How many adult tickets,a, & syudent tickets,s, were sold? Write your answer as (a,s).
Answer by rfer(16322)

a+s=480
s=480-a
------------------
8a+5s=3369
8a+5*(480-a)=3369
8a+2400-5a=3369
3a=3369-2400
3a=969
a=969/3
a=323 adults
s=480-323
s=157 students
-----------------